### v1

# The Affective Bridge: Unifying Feature Representations for Speech Deepfake Detection

Abstract
Speech deepfake detection has been widely explored using low-level acoustic descriptors. However, each study tends to select different feature sets, making it difficult to establish a unified representation for the task. Moreover, such features are not intuitive for humans to perceive, as the distinction between bona fide and synthesized speech becomes increasingly subtle with the advancement of deepfake generation techniques. Emotion, on the other hand, remains a unique human attribute that current deepfake generator struggles to fully replicate, reflecting the gap toward true artificial general intelligence. Interestingly, many existing acoustic and semantic features have implicit correlations with emotion. For instance, speech features recognized by automatic speech recognition systems often varies naturally with emotional expression. Based on this insight, we propose a novel training framework that leverages emotion as a bridge between conventional deepfake features and emotion-oriented representations. Experiments on the widely used FakeOrReal and In-the-Wild datasets demonstrate consistent and substantial improvements in accuracy, up to approximately 6% and 2% increases, respectively, and in equal error rate (EER), showing reductions of up to about 4% and 1%, respectively, while achieving comparable results on ASVspoof2019. This approach provides a unified training strategy for all features and interpretable feature direction for deepfake detection while improving model performance through emotion-informed learning.
Index Terms:
Speech, Deepfake detection, Speech, Emotion, Feature selection, Pretraining
I Introduction
Speech Deepfake detection in audio heavily relies on the selection of appropriate and robust feature representations [iqbal2022deepfake]. Numerous studies have explored different feature types to identify the most representative ones for this task. Physical acoustic features are among the most commonly used, including Mel Frequency Cepstral Coefficients (MFCCs) [hamza2022deepfake], chroma features [bakken2025deep], and comprehensive sets of acoustic descriptors extracted using tools such as openSMILE [pascu2025easy, eyben2010opensmile]. In parallel, deep learning–based features derived from pretrained models such as Whisper [radford2023robust], regarded as raw audio physical features, have also shown promising results.
Beyond these low-level representations, high-level task-oriented features have proven effective as well. These are often obtained from models trained for specific downstream applications, such as automatic speech recognition (ASR) [mansoor2024audio] and speaker verification (SV) [pianese2022deepfake], and tend to offer more interpretable representations compared to simple physical features. Despite these advances, the landscape of audio deepfake detection remains fragmented: there is no consensus on an optimal feature set. Moreover, humans often struggle to perceptually distinguish between bona fide and synthetic audio, as the generative models are largely developed and their attention is typically guided by global parts rather than fine-grained acoustic details that machine models may exploit [somoray2025human].
However, human could recognise emotion inconsistency from deepfake audio more easily compared to other features, as emotion remains a challenging aspect for machines to imitate authentically, highlighting a persistent gap between current artificial intelligence (AI) systems and general intelligence [sun2024friendlyaicomprehensivereview, mobbs2025emotion, li2025artificialemotionsurveytheories]. This limitation also underscores the interpretability of emotional processes from a human-centered viewpoint. Moreover, emotional cues have been shown to serve as effective discriminative features for audio deepfake detection on the machine side [conti2022deepfake, mittal2020emotions]. Therefore, emotion can serve as a connecting bridge between human perception and machine representation, potentially guiding the development of more interpretable and unified deepfake detection frameworks.
Emotion can also be represented through various sets of features, as demonstrated by studies utilising such features for emotion recognition. These include low-level descriptors of physical characteristics [kishore2013emotion, li2025gatedxlstmmultimodalaffectivecomputing] and higher-level, application-based feature representations such as those derived from ASR systems [tits2018asr]. The Macro-Voice framework seeks to disentangle speaker identity information from emotional cues to obtain a more purified emotion representation, thereby illustrating the intersection between SV representations and emotional features [tian2025marcovoicetechnicalreport]. This literature shows a potential that emotion representations may serve as an effective bridge for transferring feature sets to tasks, such as deepfake detection.
To address the lack of an optimal and interpretable feature representation, we propose a unified novel pre-training stage prior to the final deepfake detection classifier. This stage transfers knowledge from diverse feature domains into an emotion-based representation while mitigating catastrophic forgetting [aleixo2023catastrophic] of previously learned tasks. Although prior studies have demonstrated the effectiveness of aligning emotion features with other modalities [mittal2020emotions, hosler2021deepfakes], they primarily inject affective cues into the pipeline rather than leveraging emotion as a bridging representation between existing features. Conventional fusion strategies, such as early or late fusion [lei2025deepfake, alsaeedi2025audio], not only require careful feature selection but also often produce representations that are difficult to interpret or reuse for original tasks (e. g., ASR) or emotion recognition. In contrast, our approach preserves the physical meaning of the features, enabling a given representation to progressively learn emotion-related characteristics and transfer them effectively to deepfake detection. Experimental results show consistent improvements on the FakeOrReal (FoR) [reimao2019dataset] and InTheWild (ITW) [muller2022does] benchmarks, specifically up to approximately 6% and 2% increases for accuracy, respectively, and in equal error rate (EER), showing reductions of up to about 4% and 1%, and comparable performance on ASVSpoof2019 LA [wang2020asvspoof].
In summary, our main contributions are twofold. First, we propose a unified novel training framework that transfers and aligns heterogeneous feature sets into an emotion-centered representation, yielding a robust and interpretable feature space for deepfake detection. In this framework, both the original and emotion-related features are preserved and effectively fused, preventing catastrophic forgetting and ensuring that the original representations remain functional. Furthermore, we demonstrate that employing affective bridge: emotion as a bridging representation consistently enhances deepfake detection performance without additional training on the deepfake detection task, thereby ensuring scalability and reducing the risk of overfitting on deepfake detection. This also suggests that emotion-related features remain a distinctive cue that current deepfake audio models fail to accurately reproduce.
II Methodology: The Affective Bridge: EmoBridge
To address the aforementioned challenge of selecting appropriate features for deepfake detection, we align affective cues through a continual learning framework that jointly optimises existing features with an emotion-related task. This continual adaptation enables the feature representation to incorporate new affective information without forgetting previously learned knowledge, which is shown reliable from the continuous learning principle [de2021continual, li2025neuroplasticityartificialintelligence]. The overall framework is illustrated in Figure 1.
Figure 1: Emotion as a bridge framework. The pre-trained model for a specific task provides an original feature representation, where essential information is stored in the encoder. We further train the encoder, together with a fully connected layer, to perform an emotion recognition task, thereby fusing emotion-related features within the encoder (EmoBridge step). The outputs from each encoder layer, which now incorporate both original and emotion-based features, are then used as inputs to a classifier—such as a support vector machine (SVM)—for the final deepfake detection task.
Given the vast variety of feature representations, we consider the last hidden states of the feature encoder (for attention-based architectures) or the final layer outputs (for other deep learning architectures) as the optimal feature representations for each modality. We categorise commonly used features into three groups: (1) acoustic features, which can be extracted using openSMILE; (2) deep learning (DL) raw features, derived directly from deep neural models; and (3) application-based features, such as semantic features from ASR pre-trained models or perceptual features from SV pre-trained models.
In detail, to investigate the effect of fusing emotion representations, we select four representative feature types: acoustic features extracted with openSMILE [eyben2010opensmile]; low-level deep learning features from Whisper [radford2023robust] [^1] https://huggingface.co/openai/whisper-small ; semantic and linguistic features from the ASR model SpeechT5 [ao2022speecht5unifiedmodalencoderdecoderpretraining] [^2] https://huggingface.co/microsoft/speecht5_asr ; and perceptual features from the SV model WavLM [Chen_2022] [^3] https://huggingface.co/microsoft/wavlm-base-sv . These pre-trained models are widely adopted in their respective domains, and we maintain comparable model configurations across experiments to ensure a fair comparison.
Since our goal is to fuse emotion representations rather than optimise for the original downstream tasks, we freeze all downstream components such as decoders during training. openSMILE is treated as a special case, as it directly extracts acoustic features (e. g., MFCCs). For this model, our analysis focuses on comparing two functional-level feature sets: eGeMAPSv02, a set designed by experts and IS09, a feature set from a challenge, both are designed for emotion recognition. The latter emphasizes more emotional cues in addition to basic acoustic features. For other feature sets, we train the encoder and subsequent classifier on an emotion recognition task using a general emotion dataset. To construct a diverse and representative training corpus, we combine four widely used emotion recognition datasets: the Toronto Emotional Speech Set (TESS) [dupuis_pichora-fuller_2010_tess], Surrey Audio-Visual Expressed Emotion (SAVEE) [jackson_haq_savee], CREMA-D [cao2014crema-d], and RAVDESS [livingstone2018ravdess]. All audio from these datasets is used to train the EmoBridge stage, while we use Emotion Speech Dataset (ESD) [zhou2021esd] for testing the learned emotion representations.
After completing the EmoBridge step, we use the trained encoder to perform the final deepfake detection task. To evaluate the effectiveness of the integrated features, we extract outputs from each encoder layer and feed them individually into a SVM [cortes1995svm] for deepfake detection. To avoid feature interference in the deepfake detection process, the encoder is kept frozen during training. The average performance across all layers is then reported to quantify the contribution of our training framework. It is worth noting that our approach is scalable: if no suitable pre-trained model exists for a given feature set, a model can be trained during the pre-training stage to serve as a ready encoder. Accordingly, our emotion as a bridge framework can be expressed in Equation 1,
$\begin{split}\theta^{*}_{pre},\omega^{*}_{pre}&=\arg\max_{\theta,\omega}\log P(y_{A}\mid x_{A};\theta,\omega),\\[2.0pt] \theta^{*}_{emo+pre},\phi^{*}_{emo}&=\arg\max_{\theta,\phi}\log P(y_{B}\mid x_{B};\theta^{*}_{pre},\phi),\\[2.0pt] \varsigma^{*}_{df}&=\arg\max_{\varsigma}\log P(y_{C}\mid x_{C};\theta^{*}_{emo+pre},\varsigma),\end{split}$ (1)
where $\theta$ denotes the parameters of upstream model where feature information is stored. $\omega$, $\phi$, and $\varsigma$ are the downstream task for the pretrained task, the emotion recognition task, and the deepfake detection task, respectively. $(x_{A},y_{A})$, $(x_{B},y_{B})$, and $(x_{C},y_{C})$ are the input-output pairs for the corresponding tasks, and $\mathcal{D}_{A}$, $\mathcal{D}_{B}$, and $\mathcal{D}_{C}$ are their respective data distributions. $P(y\mid x;\cdot)$ represents the conditional probability of the label given the input, which corresponds to the predictive distribution used in cross-entropy or negative log-likelihood.
III Experiments and result
III-A Dataset and experiments setup
We select the ASVSpoof2019 LA subset [wang2020asvspoof] and the FoR [reimao2019dataset] and ITW [muller2022does] datasets as benchmarks for the deepfake detection task. For ASVSpoof2019 LA and ITW, we follow the official dataset splits provided in the original papers. For FoR, we randomly divide the dataset into training, validation, and test sets with an 8:1:1 ratio, using a random seed of 42 for both NumPy and PyTorch to ensure reproducibility. This setup results in approximately 71k, 4.5k, and 32k test samples for ASVSpoof2019 LA, FoR, and ITW, respectively.
We employ a three-layer fully connected network with hidden dimensions of 768, 256, and 7 to train the emotion recognition task. The model is optimized using AdamW with a learning rate of 1e-5 for a maximum of 40 epochs. To ensure a fair comparison, we also evaluate a pre-trained model that already incorporates emotion-related features, namely HuBERT [hsu2021hubertselfsupervisedspeechrepresentation] pre-trained on the IEMOCAP dataset [busso2008iemocap] [^4] https://huggingface.co/superb/hubert-base-superb-er. The emotion recognition results are summarised in Table I. The models achieve commendable results on the emotion test sets, indicating that the emotion fusion process is successful. Although they do not reach the state-of-the-art models for emotion recognition, this is expected, as they retain the original features and not overfited on emotion recognition.
TABLE I: EmoBridge performance. The table reports the weighted accuracy for the four pre-trained models evaluated in this study, with comparison to other models in literature.
Metrics ASR SV DL-raw Emotion I. Shahin [shahin2019emotion] S. Hamsa [hamsa2020emotion] Koya S. [koya2022ea]
Weighted Acc. 0.728 0.701 0.786 0.732 0.840 0.909 0.917
After completing the EmoBridge step, we apply the default settings of SVM from the scikit-learn library [^5] https://scikit-learn.org/stable/api/sklearn.svm.html to each layer of the trained EmoBridge models, extracting both low-level and high-level feature representations, which better represent the model feature fusion. Model performance is then evaluated using average accuracy and equal error rate (EER) as metrics.
III-B Results
For openSMILE, we selected two emotion-related feature sets, with the latter emphasizing more emotional information. For the other four models, we applied our EmoBridge strategy. Table II reports the layer-wise averaged EER and accuracy of the four models and acoustic features, evaluated on the FoR, ITW, and ASVSpoof2019 LA datasets.
TABLE II: Comparison of layer-wise averaged EER and accuracy between the Pre-trained model and the model with the emotion as a bridge strategy on three datasets.
Dataset Model Pre-trained Emotion fused
EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
FoR openSMILE .406 .574 .275 .663
Emotion .139 .854 .078 .913
SV .136 .858 .089 .908
ASR .082 .887 .045 .946
DL-raw .091 .890 .062 .931
ITW openSMILE .200 .897 .188 .897
Emotion .050 .953 .049 .953
SV .061 .942 .060 .946
ASR .054 .948 .053 .947
DL-raw .033 .969 .020 .982
ASVSpoof 2019 LA openSMILE .406 .574 .275 .663
Emotion .061 .970 .071 .966
SV .070 .964 .079 .960
ASR .076 .960 .078 .960
DL-raw .052 .946 .049 .967
This demonstrates that incorporating the emotion as a bridge strategy consistently improves the model’s discriminative capability compared to the pre-trained baseline across most datasets and feature types. Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy, indicating a trend particularly pronounced for systems using all features, where emotion-guided fusion effectively enhances the separability between genuine and fake speech. These results confirm that emotion-related feature fusion provides complementary information that benefits the deep representation for the deepfake detection task. Compared with the Emotion-pretrained HuBERT model, the performance before EmoBridge training is not dominant. We attribute this to the fact that the model was trained solely on the IEMOCAP dataset, which provides a limited and domain-specific representation of emotion. After fusing more diverse emotion knowledge through the EmoBridge strategy, the performance improves noticeably, suggesting that a broader and more balanced emotion distribution is crucial for model performance.
Moreover, across all datasets, the DL-raw features consistently benefit from the emotion as a bridge strategy, indicating that raw deep learning representations capture key cues for deepfake detection and that emotion information serves as an important complementary signal. The FoR corpus benefits the most from our strategy, showing substantial reductions in EER (e. g., from 0.082 to 0.045 for ASR features) along with notable accuracy gains. Improvements on the ITW dataset are also evident, particularly for deep-learned features, where the EER decreases from 0.033 to 0.020 and the accuracy increases from 0.969 to 0.982. These findings suggest that emotion-based bridging provides the greatest advantage in diverse or emotionally expressive conditions.
In contrast, the ASVSpoof shows comparable performance. To identify the subset where our strategy is most vulnerable, we further analyze model performance across different sources in the test set, shown in Table III.
TABLE III: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy. In this table, only the most noteworthy results are emphasised in bold.
Source ASR SV DL-raw Emotion
EmoBridge? No Yes No Yes No Yes No Yes
A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998
A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999
A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977
A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999
A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998
A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000
A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999
A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993
A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984
A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975
A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903
Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564
The results indicate that the observed degradation mainly occurs for real audio under our strategy, likely due to the imbalanced training distribution affecting the emotion-fused features. From the DL-raw results, performance on the real subset remains higher, and overall accuracy improves, suggesting that emotion-enhanced features still contribute positively. Notably, audio generated by methods A17, A18, and A19 shows more consistent improvements than other source types, as these fakes are based on real human speech, whereas other methods rely on text or TTS synthesis. This suggests that human speech provides richer prosodic and emotional cues, resulting in more realistic fakes and more detectable emotional inconsistencies, while text- or TTS-based synthesis lacks natural emotional variation, yielding smaller gains from emotion-based modeling. Overall, our strategy remains effective, and as deepfake generation advances toward more expressive synthesis, fusing emotion-related information is expected to become very important for future detection systems.
III-C Explainability for model improvement
Our approach yields performance gains, so we further investigate the model’s internal mechanism and feature representation. Specifically, we examine the mean attention values across time steps for different layers before and after applying our strategy, using the DL-raw features as an example, shown in Figure 2. The results show that the model after applying our strategy retains partial overlap with the original attention distribution, indicating that it preserves useful prior information, while also developing new attention regions where more affective cues are likely captured.
Figure 2: Comparison of mean attention values of Whisper layers before (model_ori) and after emotion as a bridge (model_new).
Additionally, the previous claim is further supported through feature visualization using selected samples from the EmoFake dataset [zhao2024emofake]. These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations. We analyze SV and ASR features, as shown in Figure 3, using t-SNE [maaten2008visualizing] for visualization.
After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion feature fusion. Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, though the clusters are slightly less compact than before EmoBridge training. This suggests that our strategy avoids catastrophic forgetting during emotion fusion and prevents overfitting to specific tasks, thereby enhancing scalability.
Figure 3: t-SNE for SV (left) and ASR (right) visualizations of selected sample representations obtained before (model_ori) and after (model_new) Emobrigde. The x_y_zx\_y\_z label denotes the speaker, emotion, and content respectively; identical subscripts indicate the same speaker, emotion, or content.
The results show that the trends for emotion and deepfake detection representations are quite similar, indicating the importance of emotion features in deepfake detection. Lower layers capture more basic features, which are less informative for application-level tasks such as emotion recognition, while higher layers provide better representations for both emotion and deepfake detection.
IV Conclusion
We proposed the Emobridge strategy, a unified approach, to incorporate emotion information into pre-trained features. This approach not only preserves the original feature representations while adding affective cues, but also improves deepfake detection performance on multiple benchmarks. We consider emotion to be a more interpretable feature, perceivable by humans, which can serve as a bridge to obtain a unified feature representation. In future work, a unified representation for deepfake detection could be achieved by exploring methods to effectively combine representations from different pre-trained models through emotion.

### v2

@@
-The Affective Bridge: Unifying Feature Representations for Speech Deepfake Detection
+Li Lv Wang Luo Zhang Schuller
 Abstract
-Speech deepfake detection has been widely explored using low-level acoustic descriptors.
-However, each study tends to select different feature sets, making it difficult to establish a unified representation for the task.
-Moreover, such features are not intuitive for humans to perceive, as the distinction between bona fide and synthesized speech becomes increasingly subtle with the advancement of deepfake generation techniques.
-Emotion, on the other hand, remains a unique human attribute that current deepfake generator struggles to fully replicate, reflecting the gap toward true artificial general intelligence.
-Interestingly, many existing acoustic and semantic features have implicit correlations with emotion.
-For instance, speech features recognized by automatic speech recognition systems often varies naturally with emotional expression.
-Based on this insight, we propose a novel training framework that leverages emotion as a bridge between conventional deepfake features and emotion-oriented representations.
-Experiments on the widely used FakeOrReal and In-the-Wild datasets demonstrate consistent and substantial improvements in accuracy, up to approximately 6% and 2% increases, respectively, and in equal error rate (EER), showing reductions of up to about 4% and 1%, respectively, while achieving comparable results on ASVspoof2019.
-This approach provides a unified training strategy for all features and interpretable feature direction for deepfake detection while improving model performance through emotion-informed learning.
-Index Terms:
-Speech, Deepfake detection, Speech, Emotion, Feature selection, Pretraining
-I Introduction
-Speech Deepfake detection in audio heavily relies on the selection of appropriate and robust feature representations [iqbal2022deepfake].
-Numerous studies have explored different feature types to identify the most representative ones for this task.
+Speech deepfake detection (DFD) has benefited from diverse acoustic and semantic speech representations, many of which encode valuable speech information and are costly to train.
+Existing approaches typically enhance DFD by tuning the representations or applying post-hoc classification on frozen features, limiting control over improving discriminative DF cues without distorting original semantics.
+We find that emotion is encoded across diverse speech features and correlates with DFD.
+Therefore, we introduce a unified, feature-agnostic, and non-destructive training framework that uses emotion as a bridging constraint to guide speech features toward DFD, treating emotion recognition as a representation alignment objective rather than an auxiliary task, while preserving the original semantic information.
+Experiments on FakeOrReal and IntheWild show accuracy improvements of up to 6% and 2%, respectively, with corresponding reductions in equal error rate.
+Code is in the supplementary material.
+keywords:
+Speech, Deepfake detection, Emotion, Feature selection, Pretraining
+1 Introduction
+Speech deepfake detection (DFD) in audio heavily relies on the selection of appropriate and robust feature representations [iqbal2022deepfake].
 Physical acoustic features are among the most commonly used, including Mel Frequency Cepstral Coefficients (MFCCs) [hamza2022deepfake], chroma features [bakken2025deep], and comprehensive sets of acoustic descriptors extracted using tools such as openSMILE [pascu2025easy, eyben2010opensmile].
@@
 Beyond these low-level representations, high-level task-oriented features have proven effective as well.
-These are often obtained from models trained for specific downstream applications, such as automatic speech recognition (ASR) [mansoor2024audio] and speaker verification (SV) [pianese2022deepfake], and tend to offer more interpretable representations compared to simple physical features.
-Despite these advances, the landscape of audio deepfake detection remains fragmented: there is no consensus on an optimal feature set.
-Moreover, humans often struggle to perceptually distinguish between bona fide and synthetic audio, as the generative models are largely developed and their attention is typically guided by global parts rather than fine-grained acoustic details that machine models may exploit [somoray2025human].
-However, human could recognise emotion inconsistency from deepfake audio more easily compared to other features, as emotion remains a challenging aspect for machines to imitate authentically, highlighting a persistent gap between current artificial intelligence (AI) systems and general intelligence [sun2024friendlyaicomprehensivereview, mobbs2025emotion, li2025artificialemotionsurveytheories].
-This limitation also underscores the interpretability of emotional processes from a human-centered viewpoint.
-Moreover, emotional cues have been shown to serve as effective discriminative features for audio deepfake detection on the machine side [conti2022deepfake, mittal2020emotions].
-Therefore, emotion can serve as a connecting bridge between human perception and machine representation, potentially guiding the development of more interpretable and unified deepfake detection frameworks.
-Emotion can also be represented through various sets of features, as demonstrated by studies utilising such features for emotion recognition.
-These include low-level descriptors of physical characteristics [kishore2013emotion, li2025gatedxlstmmultimodalaffectivecomputing] and higher-level, application-based feature representations such as those derived from ASR systems [tits2018asr].
-The Macro-Voice framework seeks to disentangle speaker identity information from emotional cues to obtain a more purified emotion representation, thereby illustrating the intersection between SV representations and emotional features [tian2025marcovoicetechnicalreport].
-This literature shows a potential that emotion representations may serve as an effective bridge for transferring feature sets to tasks, such as deepfake detection.
-To address the lack of an optimal and interpretable feature representation, we propose a unified novel pre-training stage prior to the final deepfake detection classifier.
-This stage transfers knowledge from diverse feature domains into an emotion-based representation while mitigating catastrophic forgetting [aleixo2023catastrophic] of previously learned tasks.
-Although prior studies have demonstrated the effectiveness of aligning emotion features with other modalities [mittal2020emotions, hosler2021deepfakes], they primarily inject affective cues into the pipeline rather than leveraging emotion as a bridging representation between existing features.
-Conventional fusion strategies, such as early or late fusion [lei2025deepfake, alsaeedi2025audio], not only require careful feature selection but also often produce representations that are difficult to interpret or reuse for original tasks (e.
-g., ASR) or emotion recognition.
-In contrast, our approach preserves the physical meaning of the features, enabling a given representation to progressively learn emotion-related characteristics and transfer them effectively to deepfake detection.
+These are often obtained from models trained for specific downstream applications, such as automatic speech recognition (ASR) [mansoor2024audio] and speaker verification (SV) [pianese2022deepfake], and tend to offer more representations with semantic information compared to simple physical features.
+Given the extracted features, a common approach is to train the entire network end-to-end by jointly optimising the feature extractors and the downstream classifier.
+Under this setting, pretrained models such as Whisper or wav2vec 2.0 [luo2024whisper, wang2025awaveformer] are directly fine-tuned for DFD, resulting in substantial structural adaptation of the original representations toward DFD features.
+While effective for detection, full fine-tuning substantially alters the feature space and limits the reuse of these representations in future tasks.
+Preserving their semantic information is therefore crucial, as the same features may later be required after DFD for tasks such as ASR and SV in multimodal large language models [xu2025qwen2].
+Moreover, end-to-end fine-tuning typically needs to be repeated for each new deepfake dataset, as the learned representations may overfit to dataset-specific characteristics [khanjani2023audio].
+On the other hand, to reduce training cost, an alternative post-hoc strategy directly applies a classifier head on top of frozen speech representations for DFD [saha2024exploring].
+While this approach is computationally efficient, it treats the extracted features as fixed and provides no structured mechanism to guide the representations toward deepfake-discriminative characteristics, often resulting in limited performance gains.
+Despite these advances, the field of audio DFD remains fragmented, with no unified training framework that both preserves the semantic information of the original features and guides them toward deepfake-discriminative representations.
+Notably, we observe that the emotion recognition task provides a shared factor that can guide heterogeneous speech representations toward deepfake discrimination.
+Firstly, emotion can be expressed through a variety of feature sets, including low-level descriptors of physical characteristics [kishore2013emotion, li2025gatedxlstmmultimodalaffectivecomputing] as well as higher-level application-based features derived from ASR systems [tits2018asr].
+For instance, the Macro-Voice framework disentangles speaker identity from emotional cues to obtain a more purified emotion representation, highlighting the intersection between speaker verification representations and emotional features [tian2025marcovoicetechnicalreport].
+These observations suggest that most acoustic features capture emotional information and can be used for emotion recognition.
+Moreover, emotion cues have been shown to provide effective discriminative information for audio DFD [conti2022deepfake, mittal2020emotions] and are inherently difficult for generative models to reproduce authentically over successive iterations, making them a robust and reliable signal for DFD in practical applications [sun2024friendlyaicomprehensivereview, mobbs2025emotion, li2025artificialemotionsurveytheories].
+Therefore, emotion recognition can act as a bridge for leveraging heterogeneous speech features in DFD.
+To address the lack of a unified, feature-preserving approach for audio DFD, we make the following contribution.
+We propose a training pipeline that introduces a pre-training stage prior to the final post-hoc classifier, without fine-tuning the feature extractors and requiring only a lightweight classifier for detection.
+In the pre-training stage, we found emotion, as a suitable feature-agnostic constraint applicable to arbitrary speech representations, guiding them toward deepfake-discriminative cues while preserving their original semantic structure.
+This design enables deepfake-oriented representation adjustment with lower training cost than full fine-tuning.
+Our method differs fundamentally from prior studies that demonstrate the effectiveness of emotion features in DFD by aligning emotion with other modalities [mittal2020emotions, hosler2021deepfakes].
+These works primarily focus on incorporating additional modalities, such as visual information, to improve emotion recognition, and directly employ emotion representations as input features for detection.
+Moreover, our method differs from conventional fusion strategies, such as early or late fusion, that combine emotion features with other audio representations [lei2025deepfake, alsaeedi2025audio].
+These approaches rely on explicit feature concatenation or selection, which requires careful feature engineering and exhibits limited scalability as feature diversity increases.
+In contrast, we do not treat emotion as a standalone feature nor perform explicit feature fusion; instead, we leverage emotion as a bridging constraint to guide heterogeneous representations toward deepfake discrimination without replacing their original semantic structures.
 Experimental results show consistent improvements on the FakeOrReal (FoR) [reimao2019dataset] and InTheWild (ITW) [muller2022does] benchmarks, specifically up to approximately 6% and 2% increases for accuracy, respectively, and in equal error rate (EER), showing reductions of up to about 4% and 1%, and comparable performance on ASVSpoof2019 LA [wang2020asvspoof].
-In summary, our main contributions are twofold.
-First, we propose a unified novel training framework that transfers and aligns heterogeneous feature sets into an emotion-centered representation, yielding a robust and interpretable feature space for deepfake detection.
-In this framework, both the original and emotion-related features are preserved and effectively fused, preventing catastrophic forgetting and ensuring that the original representations remain functional.
-Furthermore, we demonstrate that employing affective bridge: emotion as a bridging representation consistently enhances deepfake detection performance without additional training on the deepfake detection task, thereby ensuring scalability and reducing the risk of overfitting on deepfake detection.
-This also suggests that emotion-related features remain a distinctive cue that current deepfake audio models fail to accurately reproduce.
-II Methodology: The Affective Bridge: EmoBridge
-To address the aforementioned challenge of selecting appropriate features for deepfake detection, we align affective cues through a continual learning framework that jointly optimises existing features with an emotion-related task.
-This continual adaptation enables the feature representation to incorporate new affective information without forgetting previously learned knowledge, which is shown reliable from the continuous learning principle [de2021continual, li2025neuroplasticityartificialintelligence].
+2 Methodology: Emotion-Guided Representation Alignment (EmoBridge)
+To address the aforementioned challenge of unifying a training pipeline for preserving speech representation to enhance DFD, we align affective cues through an emotion prediction head that jointly optimises existing features with an emotion-related task.
+This emotion constraint adaptation enables the feature representation to incorporate new affective information without forgetting previously learned knowledge.
 The overall framework is illustrated in Figure 1.
-Figure 1: Emotion as a bridge framework.
-The pre-trained model for a specific task provides an original feature representation, where essential information is stored in the encoder.
-We further train the encoder, together with a fully connected layer, to perform an emotion recognition task, thereby fusing emotion-related features within the encoder (EmoBridge step).
-The outputs from each encoder layer, which now incorporate both original and emotion-based features, are then used as inputs to a classifier—such as a support vector machine (SVM)—for the final deepfake detection task.
+Figure 1: Overview of the proposed Emotion-Guided Representation Alignment (EmoBridge) framework for speech DFD.
+In Phase I, raw speech is transformed into speech representations through two parallel routes: hand-crafted acoustic feature extractors (e.g., OpenSMILE) and encoders pre-trained for other speech tasks such as ASR or speaker verification.
+For hand-crafted acoustic features, a subset of features containing richer emotion-related cues is retained.
+For deep representations, an emotion prediction head is attached to the pre-trained encoder and optimized using an emotion recognition objective, guiding the representation to incorporate emotion-related information while preserving its original task-oriented structure.
+In Phase II, the resulting emotion-aligned representations—either selected acoustic features or outputs from the trained encoder—are frozen and fed into a downstream classifier (e.g., SVM or MLP) to perform DFD, classifying input speech as real or fake.
 Given the vast variety of feature representations, we consider the last hidden states of the feature encoder (for attention-based architectures) or the final layer outputs (for other deep learning architectures) as the optimal feature representations for each modality.
 We categorise commonly used features into three groups: (1) acoustic features, which can be extracted using openSMILE; (2) deep learning (DL) raw features, derived directly from deep neural models; and (3) application-based features, such as semantic features from ASR pre-trained models or perceptual features from SV pre-trained models.
-In detail, to investigate the effect of fusing emotion representations, we select four representative feature types: acoustic features extracted with openSMILE [eyben2010opensmile]; low-level deep learning features from Whisper [radford2023robust] [^1] https://huggingface.co/openai/whisper-small ; semantic and linguistic features from the ASR model SpeechT5 [ao2022speecht5unifiedmodalencoderdecoderpretraining] [^2] https://huggingface.co/microsoft/speecht5_asr ; and perceptual features from the SV model WavLM [Chen_2022] [^3] https://huggingface.co/microsoft/wavlm-base-sv .
-These pre-trained models are widely adopted in their respective domains, and we maintain comparable model configurations across experiments to ensure a fair comparison.
-Since our goal is to fuse emotion representations rather than optimise for the original downstream tasks, we freeze all downstream components such as decoders during training.
-openSMILE is treated as a special case, as it directly extracts acoustic features (e.
-g., MFCCs).
+In detail, to investigate the effect of emotion-guided representation alignment, we select four representative types of speech features: hand-crafted acoustic features extracted using openSMILE [eyben2010opensmile]; low-level deep learning features from Whisper [radford2023robust] [^1] https://huggingface.co/openai/whisper-small ; semantic and linguistic features from the ASR model SpeechT5 [ao2022speecht5unifiedmodalencoderdecoderpretraining] [^2] https://huggingface.co/microsoft/speecht5_asr ; and perceptual features from the speaker verification model WavLM [Chen_2022] [^3] https://huggingface.co/microsoft/wavlm-base-sv.
+These pre-trained models are widely adopted in their respective domains, and we maintain comparable model configurations across experiments to ensure fair and consistent evaluation.
+For hand-crafted features, we utilise openSMILE tools to directly extract acoustic features (e.
+g., F0, MFCCs).
 For this model, our analysis focuses on comparing two functional-level feature sets: eGeMAPSv02, a set designed by experts and IS09, a feature set from a challenge, both are designed for emotion recognition.
 The latter emphasizes more emotional cues in addition to basic acoustic features.
-For other feature sets, we train the encoder and subsequent classifier on an emotion recognition task using a general emotion dataset.
+For other feature sets, we attach an emotion prediction head implemented as a multilayer perceptron (MLP) to the feature encoder and jointly optimize the encoder and the emotion head on an emotion recognition task using a general-purpose emotion dataset.
+During the EmoBridge stage, encoder parameters are updated through the emotion objective only, without retaining the original pretraining objective.
 To construct a diverse and representative training corpus, we combine four widely used emotion recognition datasets: the Toronto Emotional Speech Set (TESS) [dupuis_pichora-fuller_2010_tess], Surrey Audio-Visual Expressed Emotion (SAVEE) [jackson_haq_savee], CREMA-D [cao2014crema-d], and RAVDESS [livingstone2018ravdess].
-All audio from these datasets is used to train the EmoBridge stage, while we use Emotion Speech Dataset (ESD) [zhou2021esd] for testing the learned emotion representations.
-After completing the EmoBridge step, we use the trained encoder to perform the final deepfake detection task.
-To evaluate the effectiveness of the integrated features, we extract outputs from each encoder layer and feed them individually into a SVM [cortes1995svm] for deepfake detection.
-To avoid feature interference in the deepfake detection process, the encoder is kept frozen during training.
-The average performance across all layers is then reported to quantify the contribution of our training framework.
-It is worth noting that our approach is scalable: if no suitable pre-trained model exists for a given feature set, a model can be trained during the pre-training stage to serve as a ready encoder.
-Accordingly, our emotion as a bridge framework can be expressed in Equation 1,
-$\begin{split}\theta^{*}_{pre},\omega^{*}_{pre}&=\arg\max_{\theta,\omega}\log P(y_{A}\mid x_{A};\theta,\omega),\\[2.0pt] \theta^{*}_{emo+pre},\phi^{*}_{emo}&=\arg\max_{\theta,\phi}\log P(y_{B}\mid x_{B};\theta^{*}_{pre},\phi),\\[2.0pt] \varsigma^{*}_{df}&=\arg\max_{\varsigma}\log P(y_{C}\mid x_{C};\theta^{*}_{emo+pre},\varsigma),\end{split}$ (1)
-where $\theta$ denotes the parameters of upstream model where feature information is stored.
-$\omega$, $\phi$, and $\varsigma$ are the downstream task for the pretrained task, the emotion recognition task, and the deepfake detection task, respectively.
-$(x_{A},y_{A})$, $(x_{B},y_{B})$, and $(x_{C},y_{C})$ are the input-output pairs for the corresponding tasks, and $\mathcal{D}_{A}$, $\mathcal{D}_{B}$, and $\mathcal{D}_{C}$ are their respective data distributions.
-$P(y\mid x;\cdot)$ represents the conditional probability of the label given the input, which corresponds to the predictive distribution used in cross-entropy or negative log-likelihood.
-III Experiments and result
-III-A Dataset and experiments setup
-We select the ASVSpoof2019 LA subset [wang2020asvspoof] and the FoR [reimao2019dataset] and ITW [muller2022does] datasets as benchmarks for the deepfake detection task.
-For ASVSpoof2019 LA and ITW, we follow the official dataset splits provided in the original papers.
-For FoR, we randomly divide the dataset into training, validation, and test sets with an 8:1:1 ratio, using a random seed of 42 for both NumPy and PyTorch to ensure reproducibility.
-This setup results in approximately 71k, 4.5k, and 32k test samples for ASVSpoof2019 LA, FoR, and ITW, respectively.
+All audio from these datasets is used to train the EmoBridge stage to avoid emotion dataset bias as much as possible, while we use Emotion Speech Dataset (ESD) [zhou2021esd] for testing the learned emotion representations.
+After completing the EmoBridge stage, the emotion-aligned encoder is connected to a downstream classifier to perform the final speech DFD task.
+To evaluate the effectiveness of the emotion-guided representations, we extract the outputs from each encoder layer and feed them individually into a support vector machine (SVM) [cortes1995svm] for classification.
+During this stage, the encoder is kept frozen to avoid feature interference and to ensure that DFD performance solely reflects the quality of the learned representations.
+The average performance across all encoder layers is reported to quantify the overall contribution of the proposed training framework.
+Our approach is scalable and feature-agnostic.
+The selected experimental features span a broad range of acoustic, linguistic, and perceptual representations, and serve as representative examples to demonstrate the general applicability of the EmoBridge framework, which can be readily extended to other speech feature extractors.
+When no suitable pre-trained model is available for a given feature set, a corresponding encoder can be trained during a pre-training stage and subsequently incorporated into the proposed pipeline.
+Accordingly, the proposed framework can be formally expressed as Equation 1.
+$\begin{aligned} \theta^{*}_{pre}&=\arg\max_{\theta}\log P(y_{A}\mid x_{A};\theta),\text{ feature pretraining}\\ \theta^{*}_{emo+pre},\phi^{*}_{emo}&=\arg\max_{\theta,\phi}\log P(y_{B}\mid x_{B};\theta^{*}_{pre},\phi),\text{ emotion alignment}\\ \varsigma^{*}_{df}&=\arg\max_{\varsigma}\log P(y_{C}\mid x_{C};\theta^{*}_{emo+pre},\varsigma),\text{ DFD}\end{aligned}$
+where $\theta$ denotes the parameters of model where feature information is stored.
+$\phi$ is the parameter of the emotion prediction head, and $\varsigma$ is the parameter of of the DFD classifier.
+$(x_{A},y_{A})$, $(x_{B},y_{B})$, and $(x_{C},y_{C})$ are the input-output pairs for speech feature extraction (e.g.
+ASR, SV), emotion recognition and DFD tasks, respectively, and $\mathcal{D}_{A}$, $\mathcal{D}_{B}$, and $\mathcal{D}_{C}$ are their respective data distributions.
+$P(y\mid x;\cdot)$ represents the conditional probability of the label given the input, which corresponds to the predictive distribution used in negative log-likelihood.
+3 Experiments and result
+3.1 Dataset and experiments setup
+We select the ASVSpoof2019 LA subset [wang2020asvspoof] and the FoR [reimao2019dataset] and ITW [muller2022does] datasets as benchmarks for the DFD task.
+For ASVSpoof2019 LA and ITW, we follow the official dataset splits in the original papers.
+For FoR, we use the for-norm data splits.
+This setup results in around 71k, 4.5k, and 32k test samples for ASVSpoof2019 LA, FoR, and ITW, respectively.
 We employ a three-layer fully connected network with hidden dimensions of 768, 256, and 7 to train the emotion recognition task.
@@
 To ensure a fair comparison, we also evaluate a pre-trained model that already incorporates emotion-related features, namely HuBERT [hsu2021hubertselfsupervisedspeechrepresentation] pre-trained on the IEMOCAP dataset [busso2008iemocap] [^4] https://huggingface.co/superb/hubert-base-superb-er.
-The emotion recognition results are summarised in Table I.
-The models achieve commendable results on the emotion test sets, indicating that the emotion fusion process is successful.
+The emotion recognition results are summarised in Table 1.
+The models achieve commendable results on the emotion test sets, indicating that the emotion constraining process is successful.
 Although they do not reach the state-of-the-art models for emotion recognition, this is expected, as they retain the original features and not overfited on emotion recognition.
-TABLE I: EmoBridge performance. The table reports the weighted accuracy for the four pre-trained models evaluated in this study, with comparison to other models in literature.
+Table 1: EmoBridge performance. The table reports the weighted accuracy for the four pre-trained models evaluated in this study, with comparison to other models in literature.
 Metrics ASR SV DL-raw Emotion I. Shahin [shahin2019emotion] S. Hamsa [hamsa2020emotion] Koya S. [koya2022ea]
@@
 After completing the EmoBridge step, we apply the default settings of SVM from the scikit-learn library [^5] https://scikit-learn.org/stable/api/sklearn.svm.html to each layer of the trained EmoBridge models, extracting both low-level and high-level feature representations, which better represent the model feature fusion.
-Model performance is then evaluated using average accuracy and equal error rate (EER) as metrics.
-III-B Results
-For openSMILE, we selected two emotion-related feature sets, with the latter emphasizing more emotional information.
-For the other four models, we applied our EmoBridge strategy.
-Table II reports the layer-wise averaged EER and accuracy of the four models and acoustic features, evaluated on the FoR, ITW, and ASVSpoof2019 LA datasets.
-TABLE II: Comparison of layer-wise averaged EER and accuracy between the Pre-trained model and the model with the emotion as a bridge strategy on three datasets.
-Dataset Model Pre-trained Emotion fused
+Model performance is then evaluated using average accuracy and EER as metrics.
+3.2 Results
+Table 2 reports the layer-wise averaged EER and accuracy of the four models and acoustic features, evaluated on the FoR, ITW, and ASVSpoof2019 LA datasets.
+This evaluation protocol reflects the robustness of the representation across different abstraction levels and avoids bias to particular layers.
+Table 2: Comparison of layer-wise averaged EER and accuracy between the Pre-trained model and the model with the emotion as a bridge strategy on three datasets.
+Dataset Model Pre-trained EmoBridge
 EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
@@
 ITW openSMILE .200 .897 .188 .897
-Emotion .050 .953 .049 .953
-SV .061 .942 .060 .946
-ASR .054 .948 .053 .947
+Emotion .050 .953 .047 .953
+SV .061 .942 .058 .946
+ASR .054 .948 .052 .947
 DL-raw .033 .969 .020 .982
@@
 DL-raw .052 .946 .049 .967
-This demonstrates that incorporating the emotion as a bridge strategy consistently improves the model’s discriminative capability compared to the pre-trained baseline across most datasets and feature types.
-Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy, indicating a trend particularly pronounced for systems using all features, where emotion-guided fusion effectively enhances the separability between genuine and fake speech.
-These results confirm that emotion-related feature fusion provides complementary information that benefits the deep representation for the deepfake detection task.
+This demonstrates that incorporating the EmoBridge strategy consistently improves the model’s discriminative capability for deepfake compared to the pre-trained baseline across most datasets and feature types.
+Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy.
 Compared with the Emotion-pretrained HuBERT model, the performance before EmoBridge training is not dominant.
-We attribute this to the fact that the model was trained solely on the IEMOCAP dataset, which provides a limited and domain-specific representation of emotion.
-After fusing more diverse emotion knowledge through the EmoBridge strategy, the performance improves noticeably, suggesting that a broader and more balanced emotion distribution is crucial for model performance.
-Moreover, across all datasets, the DL-raw features consistently benefit from the emotion as a bridge strategy, indicating that raw deep learning representations capture key cues for deepfake detection and that emotion information serves as an important complementary signal.
+We attribute this to the fact that HuBERT is trained on IEMOCAP, which provides a limited coverage of emotional expressions.
+By incorporating more diverse emotion knowledge, the detection performance improves noticeably, indicating that a broader and balanced emotion distribution is necessary.
+Moreover, across all datasets, the DL-raw features consistently benefit from the EmoBridge, indicating that raw deep learning representations capture key cues for DFD and that guidance provided by emotion information is effective.
 The FoR corpus benefits the most from our strategy, showing substantial reductions in EER (e.
 g., from 0.082 to 0.045 for ASR features) along with notable accuracy gains.
-Improvements on the ITW dataset are also evident, particularly for deep-learned features, where the EER decreases from 0.033 to 0.020 and the accuracy increases from 0.969 to 0.982.
+Improvements on the ITW dataset are also observed.
+Although the relative gains are modest, the large scale of the dataset means that these gains translate into a considerable number of additional correctly classified samples.
 These findings suggest that emotion-based bridging provides the greatest advantage in diverse or emotionally expressive conditions.
-In contrast, the ASVSpoof shows comparable performance.
-To identify the subset where our strategy is most vulnerable, we further analyze model performance across different sources in the test set, shown in Table III.
-TABLE III: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy. In this table, only the most noteworthy results are emphasised in bold.
-Source ASR SV DL-raw Emotion
+In contrast, comparable performance is observed on ASVSpoof.
+To identify where our approach is most vulnerable, we further analyse model performance across different sources within the test set.
+We also compare against two existing emotion-based methods: using the static emotion feature [conti2022deepfake] (results reported directly from the original paper), and a concatenating emotion feature with others (adapted from [mittal2020emotions]), which we combine HuBERT embeddings with DL-raw features (using our same experimental configuration of the models in Section 3.1).
+The results are shown in Table 3.
+Table 3: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy.
+Column Static is from work [mittal2020emotions] (— means not provided or N/A) while Column Emotion is from our strategy using Hubert as original feature.
+In this table, only the most noteworthy results are emphasised in bold.
+Source ASR SV DL-raw Emotion Static Concatenation
 EmoBridge? No Yes No Yes No Yes No Yes
-A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998
-A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999
-A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
-A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977
-A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999
-A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998
-A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
-A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000
-A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999
-A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993
-A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984
-A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975
-A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903
-Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564
-The results indicate that the observed degradation mainly occurs for real audio under our strategy, likely due to the imbalanced training distribution affecting the emotion-fused features.
-From the DL-raw results, performance on the real subset remains higher, and overall accuracy improves, suggesting that emotion-enhanced features still contribute positively.
-Notably, audio generated by methods A17, A18, and A19 shows more consistent improvements than other source types, as these fakes are based on real human speech, whereas other methods rely on text or TTS synthesis.
-This suggests that human speech provides richer prosodic and emotional cues, resulting in more realistic fakes and more detectable emotional inconsistencies, while text- or TTS-based synthesis lacks natural emotional variation, yielding smaller gains from emotion-based modeling.
-Overall, our strategy remains effective, and as deepfake generation advances toward more expressive synthesis, fusing emotion-related information is expected to become very important for future detection systems.
-III-C Explainability for model improvement
-Our approach yields performance gains, so we further investigate the model’s internal mechanism and feature representation.
+A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998 0.948 0.994
+A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999 0.988 0.962
+A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977 0.900 0.965
+A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999 0.895 0.997
+A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998 0.890 1.000
+A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.831 1.000
+A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000 0.763 0.952
+A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999 0.927 0.954
+A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993 0.898 0.988
+A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984 0.988
+A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975 0.869
+A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903 0.905
+Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564 0.832
+The results indicate that the relatively smaller gains observed on ASVSpoof2019 LA can be attributed to the highly imbalanced distribution of bona fide and spoofed samples in this dataset, where real speech constitutes only a small portion of the training data.
+Nevertheless, the DL-raw results still achieve higher performance on the real subset and improved overall accuracy, indicating that emotion-enhanced representations remain beneficial.
+Audio generated by methods A17–A19 shows more consistent improvements than other source types, likely because these methods are from real human speech, which contains richer prosodic and emotional variation than TTS-based synthesis.
+This observation does not suggest a limitation of the proposed framework, but highlights the importance of bona fide diversity when emotion is used as an alignment signal.
+3.3 Effects of DFD classifier
+We select SVM as the main classifier for simplicity, although other classifiers could also be used.
+Previous work shows that the relative performance of different feature representations remains consistent across classifier choices [chakravarty2024lightweight], even though absolute scores may vary.
+To further validate the robustness of our approach, we conduct an ablation study on the FoR dataset using a two-layer MLP classifier (hidden dimension 512 with a 2-dimensional output), as shown in Table 4.
+The results show similar performance across feature types, demonstrating the robustness of our method.
+Table 4: EER and accuracy between the Pre-trained model and the model with the EmoBridge strategy on FoR with MLP as classifier.
+Dataset Model Pre-trained Emotion fused
+EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
+FoR openSMILE .385 .554 .265 .645
+Emotion .143 .866 .083 .915
+SV .146 .864 .087 .910
+ASR .088 .898 .066 .952
+DL-raw .095 .899 .077 .943
+3.4 Preservation of speech representation
+Additionally, we evaluated the preservation of speech representation via feature visualization using selected samples from the EmoFake [zhao2024emofake].
+These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations.
+We analyze SV and ASR features, as shown in Figure 2, using t-SNE [maaten2008visualizing] for visualization.
+After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion guidance.
+Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, although the clusters are slightly less compact than before EmoBridge training.
+Directly evaluating ASR performance via word error rate after EmoBridge is non-trivial, as ASR systems rely on a tightly coupled encoder–decoder architecture, and re-training or re-adapting the decoder would introduce additional confounding factors.
+We therefore assess representation preservation through encoder-level, task-agnostic evaluation.
+Specifically, we compare speaker verification performance in a zero-shot setting.
+Using the same settings as in [aldeneh2024remove], the models before and after EmoBridge achieve EERs of 0.313 and 0.325, respectively, on VoxCeleb [nagrani2020voxceleb], indicating no huge degradation.
+This suggests that the proposed strategy avoids fully altering the original semantic information and prevents overfiting to a specific downstream task, enhancing scalability.
+Figure 2: t-SNE for SV (left) and ASR (right) visualizations of selected sample representations obtained before (model_ori) and after (model_new) Emobrigde.
+The x_y_zx\_y\_z label denotes the speaker, emotion, and content respectively; identical subscripts indicate the same speaker, emotion, or content.
+3.5 Explainability for model improvement
+Our approach yields performance gains, so we further investigate the model’s internal mechanism.
 Specifically, we examine the mean attention values across time steps for different layers before and after applying our strategy, using the DL-raw features as an example, shown in Figure 3.
@@
 Figure 3: Comparison of mean attention values of Whisper layers before (model_ori) and after emotion as a bridge (model_new).
-Additionally, the previous claim is further supported through feature visualization using selected samples from the EmoFake dataset [zhao2024emofake].
-These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations.
-We analyze SV and ASR features, as shown in Figure 3, using t-SNE [maaten2008visualizing] for visualization.
-After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion feature fusion.
-Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, though the clusters are slightly less compact than before EmoBridge training.
-This suggests that our strategy avoids catastrophic forgetting during emotion fusion and prevents overfitting to specific tasks, thereby enhancing scalability.
-Figure 3: t-SNE for SV (left) and ASR (right) visualizations of selected sample representations obtained before (model_ori) and after (model_new) Emobrigde.
-The x_y_zx\_y\_z label denotes the speaker, emotion, and content respectively; identical subscripts indicate the same speaker, emotion, or content.
-The results show that the trends for emotion and deepfake detection representations are quite similar, indicating the importance of emotion features in deepfake detection.
-Lower layers capture more basic features, which are less informative for application-level tasks such as emotion recognition, while higher layers provide better representations for both emotion and deepfake detection.
-IV Conclusion
-We proposed the Emobridge strategy, a unified approach, to incorporate emotion information into pre-trained features.
-This approach not only preserves the original feature representations while adding affective cues, but also improves deepfake detection performance on multiple benchmarks.
-We consider emotion to be a more interpretable feature, perceivable by humans, which can serve as a bridge to obtain a unified feature representation.
-In future work, a unified representation for deepfake detection could be achieved by exploring methods to effectively combine representations from different pre-trained models through emotion.
+4 Conclusion
+We propose EmoBridge, a unified strategy that leverages emotion information as a constraint to guide pre-trained feature embeddings.
+This approach preserves the original feature representations while injecting affective cues, enabling the model to better expose discriminative patterns for DFD and consistently outperform baseline methods across multiple benchmarks.
+Our experiments show that emotion information provides the effective guidance for improving representation quality in this framework.
+Future work could explore integrating multimodal features with emotion-aware guidance for DFD.
+5 Generative AI Use Disclosure
+We only use Generative AI to check the grammar issue of our paper.